RC Meeting and Run Plan, Monday, February 18

From HallCWiki
Jump to: navigation, search

Main | Online | RC | Run Plan | Official Website | Redmine | Index | Nav

RC Meeting Information

RC Daily Meetings are at 4:30 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Current Status

  • Ran x>1 nuclear targets last night. Beam was not spectacular. Able to get about 3k (out of desired 5k) stats on each target.
  • Went to Restricted Access at 10am to continue work on the HMS Q2 power supply.
  • Target training tomorrow at 3:30pm with Jian-Ping for those who need it.

Current Run Plan

Notices in Effect for Shift Crew

  • Target Operators: When the alarm handler goes white, it is disconnected from the IOC. Log it and include carlin@jlab.org every time. Acknowledge the warning. If it is still white and not coming back, restart the alarm handler. Do not permanently silence the alarm handler.
  • Missing Reference Times during the replay: check in the 50k. Check in full replay, if able. If it happens, call RC. Guidance to solve it is here: https://logbooks.jlab.org/entry/3654584
  • HMS Q2: Do not ramp to 0 (if it reaches 10A, just turn it off). Let the magnet settle at a current for about 5 minutes before changing it. The read back current box does not work. The current listed above it is the most useful indication that it is on.
  • We have no control over polarity for HMS magnets now. They don't have meaning in the GUI, either. It is currently set for positron/proton running (Q1-, Q2+ on hall probes)
  • If rotation cannot be Reset, call the RC, then call Ops for an escorted/controlled access.
  • SHMS vacuum gauge read back in the counting house is wrong. Walter verified SHMS vacuum is currently fine.

Current Issues

Planned Accesses to the Hall

  • Continued work on HMS Q2.
  • Back flush other magnets in the Hall tomorrow during RF recovery.
  • Restricted Access during Day Shift tomorrow.
  • Re-attach wooden floor stop around 57deg (HMS side)

Planned Accelerator Configuration Changes

Special Requests